Targeting a DNA sequence for editing by the CRISPR‑Cas9 system requires that a protospacer‑adjacent motif (PAM) sequence appear just downstream of the target sequence. For S. pyogenes Cas9 this PAM is the trinucleotide NGG, where N represents any nucleotide.

Assuming a fragment of DNA includes equal amounts of A, C, G, and T bases, how many bases apart, on average, would one expect to find the NGG PAM?

how many bases average PAM distance =
Answer #1

The required PAM sequence is NGG and the given DNA fragment has an equal distribution of A, C, G and T. This means that each of these bases will have 25% or 1/4 distribution in DNA. So, for a NGG sequence, probability of 'N' = 1 as any nucleotide can occur. Probability of G = 1/4. Thus, overall probability for NGG = 1* 1/4 * 1/4 = 1/16. Thus, 2 NGG sequences will be separated on an average by 16 bases apart.